1. Find the LFSE in terms of Dq and calculate the spin-only magnetic moment (in units of Bohr-magnetons) for the following complexes:
a) cis–diaqua–cis–dichloro–cis–difluorocobaltate(II) ion
b) trans–dicyanobis(ethylenediamine)iron(III) ion
c) trans–dichlorotetrakis(triphenylphosphine)nickel(II)
d) cis–dicyanobis(oxalato)manganate(II) ion

Grade:12

1 Answers

Vikas TU
5 years ago
Ligand field stablisation enthalpies is given in Dq
D= 35ze/4a5  and
q= 2e(r4)/105        
where Ze= charge on the ligands, a=distanceof the ligands from the central ion, r4 is the mean fourth power radius of the d electrons in the central ion
 
LFSE for Co2+ (d7) = 4/5∆ₒ×10Dq/∆ₒ = 8 Dq
LFSE for Fe+2 (d6) = 2/5∆ₒ×10Dq/∆ₒ = 4 Dq
LFSE for Ni2+  (d8) = 12Dq
Formula for spin-only magnetic moment = √n(n+2) where n is the number of unpaired electron
Co2+ , n=2 , √3(3×2) = 3.87
Fe3+ , n=5 √5(5×2)= 7.07
Ni+2 n=2, =2.83
Mn2+  n=5, = 7.07